# D3PO: Preference-Based Alignment of Discrete Diffusion Models

### v1

@@ Abstract
Diffusion models (Ho et al., 2020; Song et al., 2020) have achieved state-of-the-art performance across multiple domains (Austin et al., 2021; Watson et al., 2023; Anand & Achim, 2022), with recent advancements extending their applicability to discrete data (Lou et al., 2023; Shi et al., 2024; Campbell et al., 2022; 2024). However, aligning discrete diffusion models with task-specific preferences remains challenging, particularly in scenarios where explicit reward functions are unavailable. In this work, we introduce Discrete Diffusion DPO (D3PO), the first adaptation of Direct Preference Optimization (DPO) (Rafailov et al., 2024) to discrete diffusion models formulated as continuous-time Markov chains. Our approach derives a novel loss function that directly fine-tunes the generative process using preference data while preserving fidelity to a reference distribution. We validate D3PO on a structured binary sequence generation task, demonstrating that the method effectively aligns model outputs with preferences while maintaining structural validity. Our results highlight that D3PO enables controlled fine-tuning without requiring explicit reward models, making it a practical alternative to reinforcement learning-based approaches. Future research will explore extending D3PO to more complex generative tasks, including language modeling and protein sequence generation, as well as investigating alternative noise schedules, such as uniform noising, to enhance flexibility across different applications.
@@ 1 Introduction
In this work, we introduce Discrete Diffusion DPO (D3PO), the first adaptation of DPO to discrete diffusion models. Unlike continuous diffusion models, which leverage score-matching, discrete diffusion models are formulated as Continuous-Time Markov Chains (CTMCs), requiring a different optimization framework. We derive a novel loss function that directly fine-tunes discrete diffusion models using pairwise preference data while preserving fidelity to a reference distribution.
Our key contributions are as follows. Firstly, we introduce D3PO, a DPO-based optimization framework tailored for CTMCs, enabling preference alignment in discrete diffusion models without requiring an explicit reward function. Secondly, we show that under a masking-state noising process, our preference-based objective simplifies to an intuitive closed-form expression, providing theoretical insights into its effectiveness. Thirdly, we empirically validate D3PO on a structured sequence generation task, demonstrating that it successfully aligns discrete diffusion models with preferences while maintaining distributional coherence.
@@ 3 DPO for Discrete Diffusion Models
where $q(x_{n}|x_{N})$ is the discrete time equivalent of $q_{t|1}(x_{t}|x_{1})$, and $q(x_{n+1}|x_{n},x_{N})$ is the discrete time equivalent of equation 14. We omit $c$ for compactness. By substituting the transition probability expansion for rate matrices, and taking the continuous-time limit ($N\to\infty$, $\Delta t\to 0$), the final D3PO loss for CTMCs is obtained:
$L_{\textbf{{D3PO}}{}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(x_{1}^{w},x_{1}^{l})\sim\mathcal{P},t\sim\mathcal{U}[0,1]\\ x^{w}\sim q(x_{t}|x_{1}^{w}),x^{l}\sim q(x_{t}|x_{1}^{l})\end{subarray}}\log\sigma\Bigg{[}\beta\>\mathcal{D}^{\theta}_{\text{ref}}(x_{t}^{w}|x_{1}^{w})-\beta\>\mathcal{D}_{\text{ref}}^{\theta}(x_{t}^{l}|x_{1}^{l})\Bigg{]}$ (12)
@@ 4 Preliminary Experiments
Figure 1: Results for preference-based alignment using the D3PO loss. (Left) Training loss monotonically decreases over epochs. (Center) Ratio of generated sequences corresponding to odd integers increases w.r.t. reference model. (Right) Fraction of generated sequences with valid structure remains close to 1.
To validate the effectiveness of D3PO, we conduct a small-scale experiment demonstrating how the proposed loss in Equation equation 12 enables preference alignment in a discrete diffusion model. Building on the framework of Campbell et al. (2024), we first pre-train a masking-state discrete diffusion model to generate structured binary representations of integers. Specifically, each integer $i\in\{0,\dots,N\}$ is represented as a binary sequence of length $N$, denoted as $b_{i}\in{0,1}^{N}$. The first $i$ bits are set to $1$, while the remaining bits are set to 00. The pre-trained model learns to generate valid sequences that adhere to this structured encoding rather than producing arbitrary binary strings.
@@ 5 Conclusion and Future Work
We introduce Discrete Diffusion DPO (D3PO), a novel extension of the DPO framework to diffusion models formulate as continuous-time Markov chains. Our derivation yields a computationally efficient loss function that aligns the generative sampling process with preference data while preserving fidelity to the reference distribution. Experiments on a structured binary sequence generation task confirmed that D3PO successfully biases discrete diffusion models towards preferred outputs while preserving structural validity.
@@ Appendix A Appendix Structure
The appendix is structured as follows. Appendix B discusses related work, covering advancements in discrete diffusion models, fine-tuning techniques, and preference-based optimization in diffusion models. Appendix C provides a detailed derivation of the D3PO loss for discrete diffusion models, starting from a discrete-time approximation of the CTMC formulation and extending it to the continuous-time limit. Appendix D generalizes the D3PO loss to multi-dimensional data, presenting a factorized transition model that enables tractable optimization in structured sequence generation tasks. Appendix E derives the D3PO loss for the masking noise process, adapting the framework for discrete diffusion models that use an absorbing-state corruption scheme. Appendix E.1 extends the masking noise derivation to cases with additional re-masking noise, allowing for bidirectional transitions between masked and unmasked states. Appendix E.2 provides a complexity analysis of the derived loss functions for the masking state noise process, showing that preference-based fine-tuning with D3PO is computationally efficient.
@@ Appendix B Related Work
Appendix C Full Derivation of 1-Dimensional D3PO Loss
@@
Here we provide an expression for the infinitesimal transition probability $q_{t+dt|t,1}(x_{t+dt}|x_{t},x_{1})$ in terms of the conditional rate matrix $R^{q}_{t}(x_{t},x_{t+dt}|x_{1})$ which will be useful later in the derivation of the D3PO loss.
@@ C.4 D3PO Loss
Appendix D Multi-Dimensional D3PO
In this section we adapt the D3PO loss to account for $D$-dimensional data. Consider $\bm{x}\in\{1,\cdots,S\}^{D}$ is a $D$-dimensional vector with components $x^{d}$ where $d=1,\dots,D$. We derive the DPO loss for this general case. The derivation proceeds in the same way as for the 1-dimensional case above, up to equation 21. For the $D$-dimensional case have:
@@ Appendix D Multi-Dimensional D3PO
Appendix E D3PO Loss for Masking State Models
In this section we adapt the D3PO loss for the specific case of masking noise process. In $D$ dimensions we consider independent corruption processes in each dimension, similar to the factorization assumptions made in continuous diffusion models where the forward noising processes proceed independently in each dimension.
@@ E.2 Complexity Analysis for Masking Noise Process
For the masking noise process, the derived expressions for $\mathcal{D}^{\theta,d}_{t}(\bm{x})$ in Equations equation 25 and equation 26 provide a computationally efficient way to estimate the D3PO loss function. In practice, the denoising models $p_{1\mid t}^{\theta}$ and $p_{1\mid t}^{\text{ref}}$ take as input a noisy vector $\bm{x}_{t}\in\{1,\dots,S,M\}^{D}$ and output probability vectors $p_{1\mid t}(\bm{x}_{1}\mid\bm{x}_{t})\in[0,1]^{D}$. Since the loss function requires evaluating the probability of reconstructing each dimension $x_{1}^{d}$, this can be directly accessed as the $d^{\text{th}}$ component of the model’s output.
@@
When estimating the per-example D3PO loss using a batch of size $T$ to approximate the expectation over $t\sim\mathcal{U}[0,1]$, the total number of model queries scales to $2T=O(T)$. For a dataset containing $P$ preference pairs, the overall computational complexity becomes $O(PT)$, reflecting a linear dependence on both the number of preferences and batch size. This scaling ensures that preference optimization in discrete diffusion models remains computationally efficient, making it practical for large-scale generative modeling tasks.

### v2

# Preference-Based Alignment of Discrete Diffusion Models

@@ Abstract
Diffusion models (Ho et al., 2020; Song et al., 2020) have achieved state-of-the-art performance across multiple domains (Austin et al., 2021; Watson et al., 2023; Anand & Achim, 2022), with recent advancements extending their applicability to discrete data (Lou et al., 2023; Shi et al., 2024; Campbell et al., 2022; 2024). However, aligning discrete diffusion models with task-specific preferences remains challenging, particularly in scenarios where explicit reward functions are unavailable. In this work, we introduce Discrete Diffusion DPO (D2-DPO), the first adaptation of Direct Preference Optimization (DPO) (Rafailov et al., 2024) to discrete diffusion models formulated as continuous-time Markov chains. Our approach derives a novel loss function that directly fine-tunes the generative process using preference data while preserving fidelity to a reference distribution. We validate D2-DPO on a structured binary sequence generation task, demonstrating that the method effectively aligns model outputs with preferences while maintaining structural validity. Our results highlight that D2-DPO enables controlled fine-tuning without requiring explicit reward models, making it a practical alternative to reinforcement learning-based approaches. Future research will explore extending D2-DPO to more complex generative tasks, including language modeling and protein sequence generation, as well as investigating alternative noise schedules, such as uniform noising, to enhance flexibility across different applications.
@@ 1 Introduction
In this work, we introduce Discrete Diffusion DPO (D2-DPO), the first adaptation of DPO to discrete diffusion models. Unlike continuous diffusion models, which leverage score-matching, discrete diffusion models are formulated as Continuous-Time Markov Chains (CTMCs), requiring a different optimization framework. We derive a novel loss function that directly fine-tunes discrete diffusion models using pairwise preference data while preserving fidelity to a reference distribution.
Our key contributions are as follows. Firstly, we introduce D2-DPO, a DPO-based optimization framework tailored for CTMCs, enabling preference alignment in discrete diffusion models without requiring an explicit reward function. Secondly, we show that under a masking-state noising process, our preference-based objective simplifies to an intuitive closed-form expression, providing theoretical insights into its effectiveness. Thirdly, we empirically validate D2-DPO on a structured sequence generation task, demonstrating that it successfully aligns discrete diffusion models with preferences while maintaining distributional coherence.
@@ 3 DPO for Discrete Diffusion Models
where $q(x_{n}|x_{N})$ is the discrete time equivalent of $q_{t|1}(x_{t}|x_{1})$, and $q(x_{n+1}|x_{n},x_{N})$ is the discrete time equivalent of equation 14. We omit $c$ for compactness. By substituting the transition probability expansion for rate matrices, and taking the continuous-time limit ($N\to\infty$, $\Delta t\to 0$), the final D2-DPO loss for CTMCs is obtained:
$L_{\textbf{{D2-DPO}}{}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(x_{1}^{w},x_{1}^{l})\sim\mathcal{P},t\sim\mathcal{U}[0,1]\\ x^{w}\sim q(x_{t}|x_{1}^{w}),x^{l}\sim q(x_{t}|x_{1}^{l})\end{subarray}}\log\sigma\Bigg{[}\beta\>\mathcal{D}^{\theta}_{\text{ref}}(x_{t}^{w}|x_{1}^{w})-\beta\>\mathcal{D}_{\text{ref}}^{\theta}(x_{t}^{l}|x_{1}^{l})\Bigg{]}$ (12)
@@ 4 Preliminary Experiments
Figure 1: Results for preference-based alignment using the D2-DPO loss. (Left) Training loss monotonically decreases over epochs. (Center) Ratio of generated sequences corresponding to odd integers increases w.r.t. reference model. (Right) Fraction of generated sequences with valid structure remains close to 1.
To validate the effectiveness of D2-DPO, we conduct a small-scale experiment demonstrating how the proposed loss in Equation equation 12 enables preference alignment in a discrete diffusion model. Building on the framework of Campbell et al. (2024), we first pre-train a masking-state discrete diffusion model to generate structured binary representations of integers. Specifically, each integer $i\in\{0,\dots,N\}$ is represented as a binary sequence of length $N$, denoted as $b_{i}\in{0,1}^{N}$. The first $i$ bits are set to $1$, while the remaining bits are set to 00. The pre-trained model learns to generate valid sequences that adhere to this structured encoding rather than producing arbitrary binary strings.
@@ 5 Conclusion and Future Work
We introduce Discrete Diffusion DPO (D2-DPO), a novel extension of the DPO framework to diffusion models formulate as continuous-time Markov chains. Our derivation yields a computationally efficient loss function that aligns the generative sampling process with preference data while preserving fidelity to the reference distribution. Experiments on a structured binary sequence generation task confirmed that D2-DPO successfully biases discrete diffusion models towards preferred outputs while preserving structural validity.
@@ Appendix A Appendix Structure
The appendix is structured as follows. Appendix B discusses related work, covering advancements in discrete diffusion models, fine-tuning techniques, and preference-based optimization in diffusion models. Appendix C provides a detailed derivation of the D2-DPO loss for discrete diffusion models, starting from a discrete-time approximation of the CTMC formulation and extending it to the continuous-time limit. Appendix D generalizes the D2-DPO loss to multi-dimensional data, presenting a factorized transition model that enables tractable optimization in structured sequence generation tasks. Appendix E derives the D2-DPO loss for the masking noise process, adapting the framework for discrete diffusion models that use an absorbing-state corruption scheme. Appendix E.1 extends the masking noise derivation to cases with additional re-masking noise, allowing for bidirectional transitions between masked and unmasked states. Appendix E.2 provides a complexity analysis of the derived loss functions for the masking state noise process, showing that preference-based fine-tuning with D2-DPO is computationally efficient.
@@ Appendix B Related Work
Appendix C Full Derivation of 1-Dimensional D2-DPO Loss
@@
Here we provide an expression for the infinitesimal transition probability $q_{t+dt|t,1}(x_{t+dt}|x_{t},x_{1})$ in terms of the conditional rate matrix $R^{q}_{t}(x_{t},x_{t+dt}|x_{1})$ which will be useful later in the derivation of the D2-DPO loss.
@@ C.4 D2-DPO Loss
Appendix D Multi-Dimensional D2-DPO
In this section we adapt the D2-DPO loss to account for $D$-dimensional data. Consider $\bm{x}\in\{1,\cdots,S\}^{D}$ is a $D$-dimensional vector with components $x^{d}$ where $d=1,\dots,D$. We derive the DPO loss for this general case. The derivation proceeds in the same way as for the 1-dimensional case above, up to equation 21. For the $D$-dimensional case have:
@@ Appendix D Multi-Dimensional D2-DPO
Appendix E D2-DPO Loss for Masking State Models
In this section we adapt the D2-DPO loss for the specific case of masking noise process. In $D$ dimensions we consider independent corruption processes in each dimension, similar to the factorization assumptions made in continuous diffusion models where the forward noising processes proceed independently in each dimension.
@@ E.2 Complexity Analysis for Masking Noise Process
For the masking noise process, the derived expressions for $\mathcal{D}^{\theta,d}_{t}(\bm{x})$ in Equations equation 25 and equation 26 provide a computationally efficient way to estimate the D2-DPO loss function. In practice, the denoising models $p_{1\mid t}^{\theta}$ and $p_{1\mid t}^{\text{ref}}$ take as input a noisy vector $\bm{x}_{t}\in\{1,\dots,S,M\}^{D}$ and output probability vectors $p_{1\mid t}(\bm{x}_{1}\mid\bm{x}_{t})\in[0,1]^{D}$. Since the loss function requires evaluating the probability of reconstructing each dimension $x_{1}^{d}$, this can be directly accessed as the $d^{\text{th}}$ component of the model’s output.
@@
When estimating the per-example D2-DPO loss using a batch of size $T$ to approximate the expectation over $t\sim\mathcal{U}[0,1]$, the total number of model queries scales to $2T=O(T)$. For a dataset containing $P$ preference pairs, the overall computational complexity becomes $O(PT)$, reflecting a linear dependence on both the number of preferences and batch size. This scaling ensures that preference optimization in discrete diffusion models remains computationally efficient, making it practical for large-scale generative modeling tasks.